At a certain location, the solar power per unit area reaching the Earth's surface is 210 W/m2 averaged over a 24-hour day. Suppose you live in a solar powered house whose average power requirement is 3.2 kW.

At what rate must solar energy reach your solar panels if your solar panels have a typical coversion efficiency of 13 %. (In other words, what power must your solar panels receive?) Answer in both kiloWatts and Watts. Psolar =

Psolar =

Based on the power that your solar panels must receive, calculate the area that your solar panels must have in order to maintain your average household power requirement. Apanels =

How much energy does this household consume in a typical day? E =
Answer #1

use

13% of P_solar = P_required

0.13*P_solar = 3.2 kW

P_solar = 3.2 kW/0.13

P_solar = 24.6 kW   (or) 24600 W <<<<<<<<<<<------------------Answer


now use,

intensity of solar radiation, I_solar = P_solar/A_panel

==> A_panel = P_solar/I_solar

= 24600/210

= 117 m^2 <<<<<<<---------------Answer

Energy consumed in a typical day = P_required*time

= 3.2*10^3*24*60*60

= 2.76*10^8 J (or) 76.8 kWh <<<<<<<---------Answer

since 1 kWh = 3.6*10^6 J